OpenStudy (anonymous):

which of the following is |-3-6i| a. +or- 3 root 5 b. 3 root 5 c.+or- 3i root 5 d.3 root 5i

OpenStudy (anonymous):

its b . 3 root 5 because the modulus is never negative

OpenStudy (anonymous):

the modulus of a complex number is its length (r) you get it by the good old Pythagoras theorem (-3)^2+(-6)^2=45=r^2 r=3*sqrt5

OpenStudy (anonymous):

thanks yall